OpenStudy (anonymous):

If 8p=-2X^2-100= which of the following could X possibly be?

OpenStudy (anonymous):

Answer Choices

OpenStudy (anonymous):

@mathmate @Here_to_Help15 @DullJackel09 @damonte212 @fallenangelorchid

OpenStudy (cggurumanjunath):

a

OpenStudy (anonymous):

can you explain how, @CGGURUMANJUNATH ?

OpenStudy (cggurumanjunath):

sure !

OpenStudy (anonymous):

Thank you (:

OpenStudy (cggurumanjunath):

8p = - 2x^2 - 100

OpenStudy (cggurumanjunath):

add 100 on both sides

OpenStudy (anonymous):

100+8p=-2x^2 ?

OpenStudy (cggurumanjunath):

u got it !

OpenStudy (anonymous):

Now what?

OpenStudy (cggurumanjunath):

divide by 2 on both sides

OpenStudy (anonymous):

8p+100/2=-2x^2/2

OpenStudy (anonymous):

wait wouldn't it be -2 not 2 ?

OpenStudy (cggurumanjunath):

u r right !

OpenStudy (cggurumanjunath):

100+8p / -2 = x^2

OpenStudy (cggurumanjunath):

-100-8p / 2 = x^2

OpenStudy (cggurumanjunath):

sorry c not a

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

Thank you!!

OpenStudy (cggurumanjunath):

:)

OpenStudy (anonymous):

Do you mind if I ask you a few more questions @CGGURUMANJUNATH ?

OpenStudy (cggurumanjunath):

go ahead !

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = -x^2 + 8x + 2 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 1 to x = 4 represent?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

11. Based on the graph below, what are the solutions to the equation f(x)=g(x)?

OpenStudy (anonymous):

I also have two more questions but I don't want two put them all up just in case it gets too confusing

OpenStudy (anonymous):

@CGGURUMANJUNATH ?

OpenStudy (anonymous):

soln to second question (-2.25,-7.25) and (1.5,3) read from graph intersections